OpenStudy (cggurumanjunath):

How to calculate Recurring Deposit for Rs 5000 , for 1 year @ R.O.I . (Rate of Interest) 9 % compounded quarterly ? MEDAL !

OpenStudy (alexandervonhumboldt2):

5000*(1+9/100)=

OpenStudy (cggurumanjunath):

I know it should be Rs 62984 (Final Maturity Value * 100 % correct answer); But I want to know the exact formula to Calculate this !

OpenStudy (anonymous):

What is recurring deposit??

OpenStudy (cggurumanjunath):

you have given the formula for compound interest !

OpenStudy (anonymous):

What is recurring deposit?

OpenStudy (cggurumanjunath):

it is another form like fixed deposit .

OpenStudy (anonymous):

If it's compounded quarterly, doesn't this mean divide the interest rate by four to get 0.0225 and then go:\[5000*(1+0.0225)^4\]

OpenStudy (anonymous):

It is Compound Interest no?

OpenStudy (anonymous):

I'd assume so, judging by the "compounded quarterly".

OpenStudy (anonymous):

He wants Recurring Deposit Amount not Compound Interest..

OpenStudy (anonymous):

I'm stumped on that one. I've looked up "recurring deposit" and it makes no sense to me.

OpenStudy (anonymous):

See this link then: http://www.allbankingsolutions.com/recurring-deposit-calculator-india.shtml

OpenStudy (anonymous):

In this, you deposit every month 5000rs, so total deposited amount will be 60,000rs and bank gives interest according to compounded quarterly or whatever it is..

OpenStudy (anonymous):

Oh, I think this might be starting to make more sense.

OpenStudy (anonymous):

Judging by the calculator you linked me to, the interest is compounded just before the monthly payment so in the first quarter you'd get interest on Rs10,000, which is 225. For the other three quarters you get it on Rs15,000, which is 337.50 and the final Rs5,000 monthly deposit is made just after the final interest payment at the end of the year, yielding no interest. This gives a total interest payment of Rs1232.50.

OpenStudy (cggurumanjunath):

Thank God ! , After a lot of research on internet , I finally got the answer !

OpenStudy (cggurumanjunath):

R = 5000 ; i= 9/400 n = 4 M = Rs 62984 !

OpenStudy (cggurumanjunath):

Anyways , Thanks for your help ! @CHEESE3.1416 and @waterineyes

OpenStudy (cggurumanjunath):

http://everydaycalculation.com/recurring-deposit.php